# Teaching Divisibility and Binomials with Coq

### v1

@@
Abstract: The goal of this contribution is to provide worksheets in Coq for students to learn about divisibility and binomials. These basic topics are a good case study as they are widely taught in the early academic years (or before in France). We present here our technical and pedagogical choices and the numerous exercises we developed. As expected, it required additional Coq material such as other lemmas and dedicated tactics. The worksheets are freely available and flexible in several ways.
Key-words: Worksheet, Coq, Arithmetic, Divisibility, Binomial
Enseigner divisibilité et binomiaux avec Coq
Résumé : Le but de cette contribution est de fournir des feuilles d’exercices en Coq à destination des étudiants pour l’apprentissage de la divisibilité et des coefficients binomiaux. Ces domaines élémentaires sont un bon sujet d’étude car ils sont largement enseignés durant les premières années universitaires (ou avant en France). Nous présentons nos choix techniques et pédagogiques et les nombreux exercices que nous avons développés. Sans surprise, cela a nécessité des développements Coq supplémentaires tels que de nouveaux lemmes et des tactiques dédiées. Les feuilles d’exercices sont en accès libre et d’un usage flexible.
Mots-clés : Feuille d’exercices, Coq, Arithmétique, Divisibilité, Coefficient binomial
@@ 1 Introduction
As for the exercises, they follow the spirit from Software Foundations [12] that was also used at Sorbonne Paris-Nord University [9] of a Coq file with holes. We provide exercises with comments both in French and English (as the language should not be the main difficulty for the students). Depending on the difficulty of the exercise, we may provide hints or explanations. Note that we have not yet used these on students.
We provide several levels of leeway to the teacher. First, the teacher may chose among the exercise (and their order) from the worksheets (files named WS_*.v). Second, some lemmas of the core library can be seen as an exercise, such as the committee-chair identity about binomials in Section 3.2.1. Third, the level of automation is also up to the teacher. We provide all the exercise proofs without automation such as lia but it may be allowed for more advanced proofs in order to shorten them. Last but not least, the teacher chooses the environment to be used by the students, both the IDE (possibly jsCoq) and the required modules and libraries (with a limited number of lemmas, with or without automation and so on).
@@
Divisibility and binomials are also commonly studied in the first academic year by students in mathematics and the contributions of this article range from terminale to the early academic years. Section 2 is about divisibility and Section 3 about binomials.
All the Coq material (and a useful script explained later) is freely available at the following address: https://liberabaci.gitlabpages.inria.fr/contenus/ThEdu24/ThEdu24.tar.gz that compiles with Coq versions 8.16 to 8.19. The worksheets are also available (with their proof for the demo) to be tested in a browser thanks to jsCoq [7] at:
@@ 2 Divisibility
The divisibility relation on integers and basic arithmetic in general are often taught in high school and at the beginning of mathematically oriented higher education. Without neglecting historical and practical (e.g. in cryptography) aspects, there are also important pedagogical reasons to teach arithmetic. Indeed, even a very basic course in arithmetic will expose students to many proof schemes, potent and nontrivial algorithms, and predicates of very different forms (primality, divisibility) which require the students to design careful rigorous proofs (sadly, maybe for the first time). In particular, the divisibility relation has many pedagogical benefits : working with an existentially quantified predicate, showing the power of symbolic calculus to prove unexpected facts and using one of many different tools (explicit witness, remainder computation, modular arithmetics, prime factor decomposition, and so on) to solve an exercise. It also serves as an example of a non-total order relation, with algebraic compatibility results.
@@ 2.1 State of the Art and Technical Choices
The standard library defines (in Numbers.BinNums) the type Z of integers in two steps. First, positive (nonzero) natural numbers are defined as non-empty strings of bits, starting with $1$. Then, an integer number in Z is 00 or a positive number with a sign. This means that equality needs not be redefined and that operations are quite fast.
The bulk of the theory of integers in the standard library is actually implementation independent. (Most) lemmas and conventions which are shared with natural numbers are proved in the Numbers.NatInt sub-theory following an axiomatic module-based approach. Lemmas specific to integers are proved (still in an axiomatic way) in the Numbers.Integer.Abstract sub-theory. Finally, these theories are included in the concrete module ZArith.BinInt.Z.
@@ 2.2 Computation and prime numbers
Here, the reflexivity tactic checks that 62744 mod 31 and 00 are convertible, in short, in our case, that the result of the computation of the left operand is indeed 00. This kind of translation between properties and computations is at the heart of the math-comp library [8].
@@
In practice, it proved to be as inconvenient as it is pedagogically debatable. Notice the choice to consider negative numbers non-prime which contradicts the usual mathematical conventions (as well as more advanced ideal based definitions of primality). We provide some alternative formulations, as well as a Boolean-valued function primeb, which computes:
@@
Of course, our algorithm is quite naive and pales in comparison to state of the art methods such as what can be found, for instance, in the coqprime library [13]. It is simple, probably what the students expect at this point, and can compute reasonably fast up to integers of around 8 decimal digits, and does not require to import a large development.
@@ 2.3 Tactics
Generating a large number of Coq goals may seem hard to handle for students, but it should motivate them to learn selectors such as all: or 2,6: in order to avoid repetition in the proofs.
@@ 2.4 Worksheet
We found out many exercise becomes pedagogical failures when formalized. This is the case of computational exercises such as 17 | 35^228 + 84^501: Coq computes the modulo without the students having to think about a way to do it themselves. We think one such exercise is interesting for outlining computation, but they miss how to cleverly compute $a^{b}\mod c$.
Many exercise rely on the decimal representation of number, which is out of the scope of the standard library. We therefore developed definitions and lemmas to state results such as the following ones: 91 divides any number written $abcabc$; 2 divides $n$ if and only if 2 divides the last digit of $n$; 3 divides $n$ if and only of 3 divides the sum of the digits of $n$. We chose to define our own inductive type Digits rather than using a list of digit, so that the student will not have to deal with empty digits list (for the last digit for instance).
@@ 3 Binomials
(1) $\forall a,b\in\mathbb{R},\qquad(a+b)^{n}=\sum_{k=0}^{n}{n\choose k}a^{k}b^{n-k}.$
@@
(2) $\forall n,k\in\mathbb{N},\qquad{n\choose k}\,\stackrel{{\scriptstyle\mathrm{def.}}}{{=}}\,\frac{n!}{k!(n-k)!}\quad\mbox{when }k\leqslant n,\qquad{n\choose k}\,\stackrel{{\scriptstyle\mathrm{def.}}}{{=}}\,0\quad\mbox{when }n<k,$
or its variant where the left clause (when $k\leqslant n$) is replaced with (${n\choose k}\stackrel{{\scriptstyle\mathrm{def.}}}{{=}}\prod_{i=0}^{k-1}\frac{n-i}{k-i}$. Or the additive way using sums through Pascal’s rule,
(3) $\forall n,k\in\mathbb{N},\qquad{n+1\choose k+1}\,\stackrel{{\scriptstyle\mathrm{def.}}}{{=}}\,{n\choose k}+{n\choose k+1}$
together with the initializations (${n\choose 0}\stackrel{{\scriptstyle\mathrm{def.}}}{{=}}1$, and (${0\choose k}\stackrel{{\scriptstyle\mathrm{def.}}}{{=}}0$ for $k>0$. Of course, these definitions are equivalent, and choosing one makes the others lemmas. While the second one naturally defines a natural number, the first one a priori lives in the field of rational numbers. Establishing that the division is actually exact needs a demonstration. Note that the expression using factorials is actually valid on $\mathbb{N}^{2}$, except when $(n,k)=(0,1)$, provided the use of the euclidean division and of both conventions $0!\stackrel{{\scriptstyle\mathrm{def.}}}{{=}}1$ and $n-k\stackrel{{\scriptstyle\mathrm{def.}}}{{=}}0$ when $n<k$.
Many properties of binomial coefficients involve finite summations of natural numbers, which are not always easy to deal with in a proof assistant. Note also that exercises given to students often deal with general purpose properties that may be reused in other proofs.
@@ 3.1 State of the Art
To start with, the offer provided by the Coq standard library [^4] https://coq.inria.fr/doc/V8.19.1/stdlib/. is not satisfactory, for two main reasons. First, the corpus of lemmas is extremely poor: three lemmas about factorials, five about binomial coefficients, and none about finite summations of natural numbers. Second, Pascal’s triangle exhibits in (3) an intrinsic integer nature of binomials. Thus, it is not reasonable for teaching purposes to search for assets in the Reals section of the Coq standard library. Indeed, the binomial coefficient is defined as C in Reals.Binomial, and the finite sum of endofunctions on natural numbers as sum_nat_f in Reals.Rfunctions. Moreover, binomial C is defined in $\mathbb{R}$ through the factorial formula (2) and no projection on natural numbers is provided.
@@ 3.2 Provided Library about Sums, Factorials, and Binomials
We have developed library-like modules Nat_Compl, Nat_Sum, Nat_Factorial and Nat_Binomial, and worksheet-like modules WS_Helper and WS_Binomial. The library modules contain definitions and the associated main properties. The module Nat_Compl is for complementary properties of basic operations on natural numbers. The module WS_Binomial collects proposals for exercises including statements from the library part and specific results that only have a teaching interest. The module WS_Helper aims at listing relevant results about natural numbers for solving the exercises, and at allowing their use without qualification with the Nat. prefix. It is the only module required in WS_Binomial.
The library modules are only based on PeanoNat, Compare_dec and Wf_nat from the Coq standard library. We provide definitions that are structurally identical to their usual mathematical counterparts, i.e. with increasingly indexed sums and products. There is no addition of superficial syntax, such as implicit arguments. There is no use of automation through tactics auto or lia, and only a few occurrences of easy and now. Thus, almost all proofs are completely explicit, and use plain Coq style. It is straightforward to incorporate automation in an existing proof, but the other way around is quite cumbersome, and is provided. We have also taken a particular care of the naming of the provided lemmas.
3.2.1 Definitions and Main Properties
@@
Definition sum_range a b f := if le_lt_dec a b then sum_n a (b - a + 1) f else 0.
Thus, we have sum_range 0 n f = sum_n 0 (S n). We provide extensionality results, extraction of the first and last terms, concatenation, splitting and extraction of any term, weak and strict monotony results, the zero-sum property, additivity and subtractivity, left and right distributivity of the multiplication, and left and right arbitrary index shifts.
@@
We provide Pascal’s rule, nullity and positiveness results, weak and strict monotony results, the committee-chair identity and its iterated version, equality with the factorial formula (2), the cancellation identity, rising sums, the hockey stick identity, the binomial formula,Vandermonde’s identity, and the formulas for the sum of binomials and of their square.
3.2.2 Worksheet
Note that most general properties provided in Nat_Binomial are commonly considered as good exercises for the training of students. As such, we provide a few hints and/or comments (both in French and in English) for several lemmas on binomials, that can be transmitted to students with statements stripped off their proofs. The proofs in Nat_Binomial are fully detailed, with very limited use of automation. Of course, teachers are free to let students use lia or easy, in particular if they are sufficiently advanced in the usage of Coq.
Module WS_Helper provides the lists of results from module Arith.PeanoNat.Nat in the Coq standard library that are used in each library modules described above. Thus, teachers may tune the helper module to the exercises they extract from the library modules.
Here is an example of an exercise, with annotations for the students, and solutions for the teacher, with automation (on the right) or without (on the left). Note also that the comments are provided both in French and English, so as not to introduce a potential additional difficulty. Hints for students may be the output of a Coq command, for instance:
Check binom_formula.

### v4

@@
Abstract
The goal of this contribution is to provide worksheets in Coq for students to learn about divisibility and binomials. These basic topics are a good case study as they are widely taught in the early academic years (or before in France). We present here our technical and pedagogical choices, the numerous exercises we developed and a small experiment we conducted on two students. As expected, it required additional Coq material such as other lemmas and dedicated tactics. The worksheets are freely available and flexible in several ways.
@@ 1 Introduction
As for the exercises, they follow the spirit from Software Foundations [14] that was also used at Sorbonne Paris-Nord University [11] of a Coq file with holes. We provide exercises with comments both in French and English (as the language should not be the main difficulty for the students). Depending on the difficulty of the exercise, we may provide hints or explanations.
We provide several levels of leeway to the teacher. First, the teacher may choose among the exercise (and their order) from the worksheets (files named WS_*.v). Second, some lemmas of the core library can be seen as exercises, such as the committee-chair identity about binomials in Section 3.2. Third, the level of automation is also up to the teacher. We provide all the exercise proofs without automation such as lia but it may be allowed for more advanced proofs in order to shorten them. Last but not least, the teacher chooses the environment to be used by the students, both the IDE (possibly jsCoq) and the required libraries (with a limited number of lemmas, with or without automation and so on).
@@
Divisibility and binomials are also commonly studied in the first academic year by students in mathematics and the contributions of this article range from terminale to the early academic years. Section 2 is about divisibility and Section 3 about binomials. Section 4 describes a small experiment on two students in June 2024 that has led to a several improvements of the worksheets.
All the Coq material (and a useful script explained later) is freely available at the following address:
https://liberabaci.gitlabpages.inria.fr/contenus/ThEdu24/ThEdu24.tar.gz
that compiles with Coq versions 8.16 to 8.20. The worksheets are also available (with their proof for the demo) to be tested in a browser thanks to jsCoq [9] at:
@@ 2 Divisibility
The divisibility relation on integers and basic arithmetic in general are often taught in high school and at the beginning of mathematically oriented higher education. Without neglecting historical (arithmetic was, along with geometry, one of the main mathematical sciences in the Greek antiquity) and practical (e.g. in cryptography) aspects, there are also important pedagogical reasons to teach arithmetic. Indeed, even a very basic course in arithmetic will expose students to many proof schemes, potent and nontrivial algorithms, and predicates of very different forms (primality, divisibility) which require the students to design careful rigorous proofs (sadly, maybe for the first time). In particular, the divisibility relation has many pedagogical benefits : working with an existentially quantified predicate, showing the power of symbolic calculus to prove unexpected facts and using one of many different tools (explicit witness, remainder computation, modular arithmetics, prime factor decomposition, and so on) to solve an exercise. It also serves as an example of a non-total order relation, with algebraic compatibility results.
@@ 2.1 State of the Art and Technical Choices
The standard library defines (in Numbers.BinNums) the type Z of integers in two steps. First, positive (nonzero) natural numbers are defined as non-empty strings of bits, starting with $1$. Then, an integer number in Z is 00 or a positive number with a sign. This representation of integers has many qualities, the first being that equality needs not be redefined (there is no need to use a quotient as is the case for instance with the usual mathematical construction of $\mathbb{Z}$ as equivalence classes of couples of natural numbers). The other one, maybe not relevant for an introductory mathematics course, is that this makes operations quite fast.
The bulk of the theory of integers in the standard library is actually implementation independent. (Most) lemmas and conventions which are shared with natural numbers are proved in the Numbers.NatInt sub-theory following an axiomatic module-based approach. Lemmas specific to integers are proved (still in an axiomatic way) in the Numbers.Integer.Abstract sub-theory. Finally, these theories are included in the concrete module ZArith.BinInt.Z. This axiomatic module-based approach makes it possible to factor many lemmas which are common between natural numbers and integers and also transfer for free all the theory, say of natural numbers, to any concrete implementation (say Peano natural numbers or binary natural numbers).
@@ 2.2 Computation and prime numbers
The usual notation for divisibility allows for more concise statements:
Infix "|" := divide (at level 0).
@@
Here, the reflexivity tactic checks that 62744 mod 31 and 00 are convertible, in short, in our case, that the result of the computation of the left operand is indeed 00. This kind of translation between properties and computations is the basis of the “Small Scale Reflection” methodology used in the math-comp library [10]: mathematical propositions having two sides, one computational and one logical, with reflection lemmas (akin to Z.mod_div here, but more elaborate) to toggle between those.
@@
In practice, it proved to be as inconvenient as it is pedagogically debatable. Notice the choice to consider negative numbers non-prime which contradicts more advanced ideal based definitions of primality, but is common in introductory mathematical courses on arithmetic. We provide some alternative formulations, as well as a Boolean-valued function primeb, which computes:
@@
Of course, our algorithm is quite naive and pales in comparison to state-of-the-art methods such as what can be found, for instance, in the coqprime library [16]. It is simple, probably what the students expect at this point, and can compute reasonably fast up to integers of around 8 decimal digits, and does not require to import a large development.
@@ 2.3 Tactics
Generating a large number of Coq goals may seem hard to handle for students, but it should motivate them to learn selectors such as all: or 1,2: in order to avoid repetition in the proofs.
study x between 0 and 8 as h3.
10: apply Z.divide_pos_le in h2; lia.
all: subst x.
1: now apply Z.divide_0_l in h2.
1, 2, 4, 8: now repeat constructor.
Listing 1: An example from exercise15
@@ 2.4 Worksheet
We found out exercises become pedagogical failures when formalized. This is the case of computational exercises such as 17 | 35^228 + 84^501: Coq computes the modulo without the students having to think about a way to do it themselves. We think one such exercise is interesting for outlining computation, but they miss how to cleverly compute $a^{b}\mod c$.
Many exercises rely on the decimal representation of number, which is out of the scope of the standard library. We therefore developed definitions and lemmas to state results such as the following ones: 91 divides any number written $abcabc$; 2 divides $n$ if and only if 2 divides the last digit of $n$; 3 divides $n$ if and only of 3 divides the sum of the digits of $n$. We chose to define our own inductive type Digits rather than using a list of digits, so that the student will not have to deal with empty digits list (for the last digit for instance). We explored the option of only allowing a canonical representation of decimal numbers by forbidding trailing zeroes. This proved to be a bad idea in practice, so we went back to a simpler representation. Another consideration was choosing whether our list of digits was little or big endian. Both approaches have strengths and weaknesses; it is easier to convert a little endian digit list into a number :
Fixpoint Z_of_Digits_alt (l : Digits): Z :=
match l with
[…]
end.
Listing 2: Little endian
Fixpoint Z_of_Digits (l : Digits): Z :=
match l with
[…]
(Z_of_Digit x)*10^(digits_len xs) +
(Z_of_Digits xs)
end.
Listing 3: Big endian
The big endian function is both more computationally expensive and conceptually more complex than the little endian one. Despite that, we choose the big endian representation, because it is more natural to read numbers left-to-right, and the proofs are not much more difficult.
[…]
@@ 3 Binomials
After a mathematical introduction, Section 3.1 presents a (large) state of the art about binomials. Section 3.2 describes numerous additional needed lemmas, especially on sums and factorials, before going into the details of the worksheet in Section 3.3.
@@
$\forall a,b\in\mathbb{R},\qquad(a+b)^{n}=\sum_{\_}{k=0}^{n}{n\choose k}a^{k}b^{n-k}.$
@@
$\forall n,k\in\mathbb{N},\qquad{n\choose k}\,\stackrel{{\scriptstyle\mathrm{def.}}}{{=}}\,\frac{n!}{k!(n-k)!}\quad\mbox{when }k\leqslant n,\qquad{n\choose k}\,\stackrel{{\scriptstyle\mathrm{def.}}}{{=}}\,0\quad\mbox{when }n<k,$ (1)
or its variant where the left clause (when $k\leqslant n$) is replaced with (${n\choose k}\stackrel{{\scriptstyle\mathrm{def.}}}{{=}}\prod_{\_}{i=0}^{k-1}\frac{n-i}{k-i}$. Or the additive way using sums through Pascal’s rule,
$\forall n,k\in\mathbb{N},\qquad{n+1\choose k+1}\,\stackrel{{\scriptstyle\mathrm{def.}}}{{=}}\,{n\choose k}+{n\choose k+1},$ (2)
together with the initializations (${n\choose 0}\stackrel{{\scriptstyle\mathrm{def.}}}{{=}}1$ for $n\in\mathbb{N}$, and (${0\choose k}\stackrel{{\scriptstyle\mathrm{def.}}}{{=}}0$ for $k>0$. Of course, these definitions are equivalent, and choosing one makes the other a lemma. While the second one naturally defines a natural number, the first one a priori lives in the field of rational numbers. Establishing that the division is actually exact needs a demonstration. Note that the expression using factorials is actually valid on $\mathbb{N}^{2}$, except when $(n,k)=(0,1)$, provided the use of the Euclidean division and of both conventions $0!\stackrel{{\scriptstyle\mathrm{def.}}}{{=}}1$ and $n-k\stackrel{{\scriptstyle\mathrm{def.}}}{{=}}0$ when $n<k$.
Many properties of binomial coefficients involve finite summations of natural numbers, which are not always easy to deal with in a proof assistant. Note also that exercises given to students often deal with general purpose properties that may be reused in other proofs. Then, teachers would be free to give students a truncated library file and ask them to prove the missing results.
@@ 3.1 State of the Art
To start with, the offer provided by the Coq standard library [^4] https://coq.inria.fr/doc/V8.19.1/stdlib/. is not satisfactory, for three main reasons. First, the corpus of lemmas is extremely poor: three lemmas about factorials, five about binomial coefficients, and none about finite summations of natural numbers. Second, Pascal’s triangle exhibits in (2) an intrinsic integer nature of binomials. Thus, it is not reasonable for teaching purposes to search for assets in the Reals section of the Coq standard library. Indeed, the binomial coefficient is defined as C in Reals.Binomial, and the finite sum of endofunctions on natural numbers as sum_nat_f in Reals.Rfunctions. Moreover, binomial C is defined in $\mathbb{R}$ through the factorial formula (1) and no projection on natural numbers is provided. And third, some lemmas could have more friendly names: lt_0_fact could be renamed fact_gt_0, as it is a good policy to take the main notion as prefix, and pascal_step{1,2,3} should have more explicit names related to their semantics.
@@ 3.2 Provided Library about Sums, Factorials, and Binomials
We have developed library-like modules Nat_Compl, Nat_Sum, Nat_Factorial and Nat_Binomial, and worksheet-like modules WS_Helper and WS_Binomial. The library modules contain definitions and the associated main properties. The module Nat_Compl is for complementary properties of basic operations on natural numbers. The module WS_Binomial collects proposals for exercises and their possible solution, including statements from the library part and specific results that only have a teaching interest. The module WS_Helper aims at listing relevant results about natural numbers for solving the exercises, and at allowing their use without qualification with the Nat. prefix. It is the only module required in WS_Binomial.
The library modules are only based on PeanoNat, Compare_dec and Wf_nat from the Coq standard library. We provide definitions that are structurally identical to their usual mathematical counterparts, i.e. with increasingly indexed sums and products. There is no addition of superficial syntax, such as implicit arguments. There is no use of automation through tactics auto, lia, now or easy. Thus, all proofs are completely explicit, and use plain Coq style. It is straightforward to incorporate automation in an existing proof, but the other way around is quite cumbersome, and is provided.
We have also taken a particular care in naming the provided lemmas. In module Nat_Compl, the names of lemmas are prefixed by nat_, as if they come from module Arith.PeanoNat.Nat from the Coq standard library. In the other library modules, we use the following naming rules: the main notion of the lemma is always used as prefix (here sum_n_ and sum_range_ in Nat_Sum, fact_ in Nat_Factorial and binom_ in Nat_Binomial), results with reverse implication have suffix _rev, equivalence results have suffix _equiv, contrapositive version has suffix _contra, alternate version has suffix _alt, significant ingredients of the conclusion are used to fill the names, and finally, we use common names when they exists.
@@
Definition sum_range a b f := sum_n a (S b - a) f.
Thus, we have sum_range 0 n f = sum_n 0 (S n). We provide extensionality results, extraction of the first and last terms, concatenation, splitting and extraction of any term, weak and strict monotony results, the zero-sum property, additivity and subtractivity, left and right distributivity of the multiplication, left and right arbitrary index shifts, and a double summation result for the multiplication of polynomials.
@@
Among other basic results, we provide Pascal’s rule, nullity and positiveness results, weak and strict monotony results, the committee-chair identity (3) and its iterated version (4), equality with the factorial formula (1), the cancellation identity (5), rising sums (6), the hockey stick identity (7), the binomial formula (8), Vandermonde’s identity (9), and the formulas for the sum of binomials (10) and of their square (11).
$\forall n,k\in\mathbb{N},\qquad$ $(k+1){n\choose k+1}=n{n-1\choose k},$ (3)
$\forall n,k,i\in\mathbb{N},\qquad$ $(k\not=0\lor i\leqslant n)\quad\Longrightarrow\quad(k+i)!(n-i)!{n\choose k+i}=n!k!{n-i\choose k},$ (4)
$\forall n,k,i\in\mathbb{N},\qquad$ (${n\choose k+i}{k+i\choose i}={n\choose i}{n-i\choose k},$ (5)
$\forall m,n\in\mathbb{N},\qquad$$\sum_{\_}{i=0}^{n}{m+i\choose m}={m+n+1\choose m+1},$ (6)
$\forall n,k\in\mathbb{N},\qquad$$\sum_{\_}{i=k}^{n}{i\choose k}={n+1\choose k+1},$ (7)
$\forall a,b,n\in\mathbb{N},\qquad$ $(a+b)^{n}=\sum_{\_}{i=0}^{n}{n\choose i}a^{i}b^{n-i},$ (8)
$\forall m,n,p\in\mathbb{N},\qquad$ (${m+n\choose p}=\sum_{\_}{i=0}^{p}{m\choose i}{n\choose p-i},$ (9)
$\forall n\in\mathbb{N},\qquad$$\sum_{\_}{k=0}^{n}{n\choose k}=2^{n},$ (10)
$\forall n\in\mathbb{N},\qquad$$\sum_{\_}{k=0}^{n}{n\choose k}^{2}={2*n\choose n}.$ (11)
The mathematical proofs of these results do not deserve to be given here. However, basic hints are written as comments in Nat_Binomial, and more detailed guidelines are provided in the worksheet WS_Binomial, as presented in the next section.
3.3 Worksheet
Note that most general properties provided in Nat_Binomial are commonly considered as good exercises for the training of students. As such, we provide a few hints and/or comments (both in French and in English) for several lemmas on binomials, that can be transmitted to students with statements stripped of their proofs. The proofs in Nat_Binomial are fully detailed, with no use of automation. Of course, teachers are free to let students use different levels of automation, in particular if they are sufficiently advanced in the usage of Coq.
The module WS_Helper provides a sufficient corpus of results from the standard library both to reproduce the proofs of the provided library (see Section 3.2) as well as to solve the proposed exercises. Thus, it is the only module required in the worksheet WS_Binomial. Teachers may tune the helper module to the exercises they extract from the library modules. They can also choose to hide these lemmas when students use the Search command by modifying the Add Search Blacklist command. We can notice that the corpus mentioned above actually behaves as a whitelist (see the second example binom_sum_sqr below).
The worksheet WS_Binomial includes several exercises (lemmas) with their possible solution (proofs). They are equipped with comments to help students. Comments are provided in several languages (here in English and in French), so as not to introduce a potential additional difficulty. Comments may be provided with several levels of mathematical details. A script automatically generates the student version with the selected options. It replaces the tagged solution with the command Admitted so that the file is still compiling. It is also capable of outputting HTML using jsCoq, turning the comments into rich text using the coqdoc [7] format.
We detail now three examples from WS_Binomial. The first one, binom_ex01a and its variant binom_ex01a_89, illustrate on a binomial manipulation exercise the possibility to give students more or less precise indications, tagged here fr1/en1 for a precise indication, and fr2/en2 for more advanced students (not yet handled by the script).
(*fr L’exercice suivant provient de *) (*en The following exercise is taken from *)
(* Hyperbole Terminale - Spécialité Mathématiques, Nathan, 2020 (in French). *)
(*fr Faire la preuve du lemme suivant et remplacer "Admitted" par "Qed". *)
(*en Fill the proof of next lemma and replace "Admitted" by "Qed". *)
Lemma binom_ex01a_89 : Lemma binom_ex01a :